# Conformal Validity Guarantees Exist for Any Data Distribution

### v2

@@ B.2 Deriving Practical Conformal Algorithms under Agent-Induced MFCS
With the conformal coverage validity guarantee for MFCS in Corollary B.1 at hand, we now turn to deriving practical algorithms for computing or estimating the conformal prediction set defined in Eq. (15) under ML-agent-induced MFCS. As we discuss in Section 4.3 of our main paper, the primary practical bottleneck in this case is computational complexity: While the epistemic challenge is overcome when MFCS is agent-induced (due to $\mathbb{P}(X_{j}\mid Z_{1},...,Z_{j-1})$ representing an ML-agent-controlled query function at time $j$), the complexity for computing the weights in Eq. (14) (equivalent to our main paper’s Eq. (4.3)) is still $\mathcal{O}(\prod_{j=1}^{t}(n+j))$, which quickly becomes intractable for large $t$. To alleviate this bottleneck, we thus proposed an estimation that only uses the “highest-order” terms from the $d$ most recent timesteps,
@@
$\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t+1-d}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t+1-d}}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})},$
where we call $d$ our “estimation depth” and $\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}$ is our “$d$-step” approximation of the MFCS weights. This $d$-step approximation of the MFCS thus has reduced complexity $\mathcal{O}(\prod_{j=t+1-d}^{t}(n+j))$, which is a tractable polynomial when $d$ is small.
@@
When the ML model controlling the query function $\mathbb{P}(X_{j}\mid Z_{1},...,Z_{j-1})$ treats its training data $Z_{1},...,Z_{j-1}$ symmetrically at each timestep (as in our experiments), we can factorize this even further and derive a recursive algorithm for computing the weights, wherein $d$ corresponds to the recursion depth. To provide intuition for this derivation and recursive implementation, let us start by letting $d=3$ and return to the simpler notation used in the main text, which drops the random variables (i.e., leaves them implicit to be more concise):
@@
$\widehat{p}_{n+t}^{\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3})}\{z_{i}\mid E_{z}^{(t)}\}=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t-2}}\mathbb{P}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t-2}}^{n+t}\mathbb{P}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{\sigma}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
@@
$\widehat{p}_{n+t}^{\ (3)}\{z_{i}\mid E_{z}^{(t)}\}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{i}\sum_{\sigma:\sigma(n+t)=i}}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
as we will now make use of knowing that $\sigma(n+t)=i$ within each of the summations $\sum_{\sigma:\sigma(n+t)=i}$ in the numerator and denominator. To do so, we need slightly more notation: for any subset of the data indices $K\subseteq\{1,...,n+t\}=[n+t]$, we write $z_{-K}=\{z_{i}:i\in[n+t]\backslash K\}$ to denote the set of data whose indices are not in $K$; for example, $z_{-\{i,j\}}=\{1,...,n+t\}\backslash\{i,j\}$. And, since we are assuming ML models that treat their data symmetrically at each timestep, we allow our conditional probabilities to ignore the ordering of data observations being conditioned on (e.g., $\mathbb{P}(x\mid z_{-K})$ does not depend on the ordering of the indices in $z_{-K}$). So, with this assumption and notation, the fact that $\sigma(n+t)=i$ within the summation $\sum_{\sigma:\sigma(n+t)=i}$ implies $\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})=\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})$. So, we can make this substitution and factor out $\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})$ from $\sum_{\sigma:\sigma(n+t)=i}$, since it no longer depends on $\sigma$:
@@
$=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\sum_{\sigma:\sigma(n+t)=i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}.$
@@
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
and similarly as before, within the summation $\sigma(n+t-1)=j$, so $\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})=\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})$. Making this substitution and factoring, and then repeating the pattern again for the third term ultimately gives
@@
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{\sigma:\sigma(n+t-1)=j}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{\sigma:\sigma(n+t-1)=j}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{k\neq i,j}\sum_{\sigma:\sigma(n+t-2)=k}}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{k\neq i,j}\sum_{\sigma:\sigma(n+t-2)=k}}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{k\neq i,j}\mathbb{P}(x_{k}\ |\ z_{-\{i,j,k\}})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{k\neq i,j}\mathbb{P}(x_{k}\ |\ z_{-\{i,j,k\}})}.$
@@
$\widehat{p}_{n+t}^{\ ({d})}\{z_{i_{1}}\mid E_{z}^{(t)}\}=\frac{\overbrace{\mathbb{P}({x_{i_{1}}}\ |\ {z_{-\{i_{1}\}}})}^{\text{recursion depth 1}}\overbrace{\textstyle\sum_{i_{2}\neq i_{1}}{\mathbb{P}(x_{i_{2}}\ |\ z_{-\{i_{1},i_{2}\}})}}^{\text{recursion depth 2}}\cdots\overbrace{\textstyle\sum_{i_{d}\neq i_{1},...,i_{d-1}}\mathbb{P}(x_{i_{d}}\ |\ z_{-\{i_{1},i_{2},...,i_{d}\}})}^{\text{recursion depth }d}}{\sum_{i_{1}}\mathbb{P}({x_{i_{1}}}\ |\ {z_{-\{i_{1}\}}})\sum_{i_{2}\neq i_{1}}{\mathbb{P}(x_{i_{2}}\ |\ z_{-\{i_{1},i_{2}\}})}\cdots\sum_{i_{d}\neq i_{1},...,i_{d-1}}\mathbb{P}(x_{i_{d}}\ |\ z_{-\{i_{1},i_{2},...,i_{d}\}})},$

### v3

@@ B.2 Deriving Practical Conformal Algorithms under Agent-Induced MFCS
With the conformal coverage validity guarantee for MFCS in Corollary B.1 at hand, we now turn to deriving practical algorithms for computing or estimating the conformal prediction set defined in Eq. (15) under ML-agent-induced MFCS. As we discuss in Section 4.3 of our main paper, the primary practical bottleneck in this case is computational complexity: While the epistemic challenge is overcome when MFCS is agent-induced (due to $p(x_{j}\mid Z_{1},...,Z_{j-1})$ representing an ML-agent-controlled query function at time $j$), the complexity for computing the weights in Eq. (14) (equivalent to our main paper’s Eq. (4.3)) is still $\mathcal{O}(\prod_{j=1}^{t}(n+j))$, which quickly becomes intractable for large $t$. To alleviate this bottleneck, we thus proposed an estimation that only uses the “highest-order” terms from the $d$ most recent timesteps,
@@
$\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t+1-d}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t+1-d}}^{n+t}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$
where we call $d$ our “estimation depth” and $\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}$ is our “$d$-step” approximation of the MFCS weights. [^16] We note that, due to assuming the first $n$ points are IID in MFCS it follows that $p_{X_{j}|Z_{1},...,Z_{j-1}}(x\mid Z_{1},...,Z_{j-1})=p_{X}(x)$ for $j\in\{1,...,n\}$, which can allow for including the terms corresponding to $j\in\{1,...,n\}$ in the product with only marginally more computation, regardless of $d$ (whereas as we have defined $\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}$, selecting $d\in\{1,...,t\}$ excludes these terms from the product). However, in certain cases such as uniform random IID initialization where all points initial density on each point is a constant, these terms cancel out in the ratio and thus can be ignored without cost. Accordingly, for this reason and for ease of exposition, we focus on the estimated weights $\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}$ as we have introduced above. This $d$-step approximation of the MFCS thus has reduced complexity $\mathcal{O}(\prod_{j=t+1-d}^{t}(n+j))$, which is a tractable polynomial when $d$ is small. This estimation depth $d$ can be specified by a user based on their computational budget, with larger values of $d$ better approximating the coverage validity guarantee at the expense of increasing computational demand.
@@
When the ML model controlling the query function $p(x_{j}\mid Z_{1},...,Z_{j-1})$ treats its training data $Z_{1},...,Z_{j-1}$ symmetrically at each timestep (as in our experiments), we can factorize this even further and derive a recursive algorithm for computing the weights, wherein $d$ corresponds to the recursion depth. To provide intuition for this derivation and recursive implementation, let us start by letting $d=3$ and return to the simpler notation used in the main text, which drops the random variables (i.e., leaves them implicit to be more concise):
@@
$\widehat{p}_{n+t}^{\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3})}\{z_{i}\mid E_{z}^{(t)}\}=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t-2}}p(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t-2}}^{n+t}p(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}p(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{\sigma}p(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
@@
$\widehat{p}_{n+t}^{\ (3)}\{z_{i}\mid E_{z}^{(t)}\}=$
$\frac{\sum_{\sigma:\sigma(n+t)=i}p(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{i}\sum_{\sigma:\sigma(n+t)=i}}p(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
as we will now make use of knowing that $\sigma(n+t)=i$ within each of the summations $\sum_{\sigma:\sigma(n+t)=i}$ in the numerator and denominator. To do so, we need slightly more notation: for any subset of the data indices $K\subseteq\{1,...,n+t\}=[n+t]$, we write $z_{-K}=\{z_{i}:i\in[n+t]\backslash K\}$ to denote the set of data whose indices are not in $K$; for example, $z_{-\{i,j\}}=\{1,...,n+t\}\backslash\{i,j\}$. And, since we are assuming ML models that treat their data symmetrically at each timestep, we allow our conditional probabilities to ignore the ordering of data observations being conditioned on (e.g., $p(x\mid z_{-K})$ does not depend on the ordering of the indices in $z_{-K}$). So, with this assumption and notation, the fact that $\sigma(n+t)=i$ within the summation $\sum_{\sigma:\sigma(n+t)=i}$ implies $p(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})=p({x_{i}}\ |\ {z_{-\{i\}}})$. So, we can make this substitution and factor out $p({x_{i}}\ |\ {z_{-\{i\}}})$ from $\sum_{\sigma:\sigma(n+t)=i}$, since it no longer depends on $\sigma$:
@@
$=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p({x_{i}}\ |\ {z_{-\{i\}}})}\cdot p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\sum_{\sigma:\sigma(n+t)=i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p({x_{i}}\ |\ {z_{-\{i\}}})}\cdot p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p({x_{i}}\ |\ {z_{-\{i\}}})}\sum_{\sigma:\sigma(n+t)=i}p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p({x_{i}}\ |\ {z_{-\{i\}}})}\sum_{\sigma:\sigma(n+t)=i}p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}.$
@@
$=$ $\frac{p({x_{i}}\ |\ {z_{-\{i\}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}}p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}p({x_{i}}\ |\ {z_{-\{i\}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}}p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
and similarly as before, within the summation $\sigma(n+t-1)=j$, so $p(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})=p(x_{j}\ |\ z_{-\{i,j\}})$. Making this substitution and factoring, and then repeating the pattern again for the third term ultimately gives
@@
$=$ $\frac{p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p(x_{j}\ |\ z_{-\{i,j\}})}\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p(x_{j}\ |\ z_{-\{i,j\}})}\cdot p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{\sigma:\sigma(n+t-1)=j}p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{\sigma:\sigma(n+t-1)=j}p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{p(x_{j}\ |\ z_{-\{i,j\}})}\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{k\neq i,j}\sum_{\sigma:\sigma(n+t-2)=k}}p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{p(x_{j}\ |\ z_{-\{i,j\}})}\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{k\neq i,j}\sum_{\sigma:\sigma(n+t-2)=k}}p(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{p(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{k\neq i,j}p(x_{k}\ |\ z_{-\{i,j,k\}})}{\sum_{i}p({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{p(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{k\neq i,j}p(x_{k}\ |\ z_{-\{i,j,k\}})}.$
@@
$\widehat{p}_{n+t}^{\ ({d})}\{z_{i_{1}}\mid E_{z}^{(t)}\}=\frac{\overbrace{p({x_{i_{1}}}\ |\ {z_{-\{i_{1}\}}})}^{\text{recursion depth 1}}\overbrace{\textstyle\sum_{i_{2}\neq i_{1}}{p(x_{i_{2}}\ |\ z_{-\{i_{1},i_{2}\}})}}^{\text{recursion depth 2}}\cdots\overbrace{\textstyle\sum_{i_{d}\neq i_{1},...,i_{d-1}}p(x_{i_{d}}\ |\ z_{-\{i_{1},i_{2},...,i_{d}\}})}^{\text{recursion depth }d}}{\sum_{i_{1}}p({x_{i_{1}}}\ |\ {z_{-\{i_{1}\}}})\sum_{i_{2}\neq i_{1}}{p(x_{i_{2}}\ |\ z_{-\{i_{1},i_{2}\}})}\cdots\sum_{i_{d}\neq i_{1},...,i_{d-1}}p(x_{i_{d}}\ |\ z_{-\{i_{1},i_{2},...,i_{d}\}})},$ (16)